OpenStudy (90890908967777878888):

compare the pair of numbers.using<,=,or > 3/4 33/40 and ill give you an mendal

OpenStudy (anonymous):

Which is greater? Or are they equal?

OpenStudy (anonymous):

hint \[\frac{3}{4}=\frac{30}{40}\]

OpenStudy (anonymous):

3/4 = 0.75 and 33/40 = 0.825

OpenStudy (90890908967777878888):

3/40 is grater

OpenStudy (anonymous):

You can use those decimals instead to make it easier

OpenStudy (anonymous):

3/40 isn't an option.

OpenStudy (anonymous):

You use either < or > with the mouth opening towards the greater one.

OpenStudy (90890908967777878888):

<

OpenStudy (anonymous):

so it would be 0.75 __ 0.825, in our case you can even multiply both decimals by 100 to make it a whole number

OpenStudy (anonymous):

< is correct.

OpenStudy (anonymous):

The simplest form would be 75 __ 82.5

OpenStudy (anonymous):

So yes, Ephemera is correct

OpenStudy (90890908967777878888):

is it my answer correct

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Good Job, hope that helped!